I am not even sure how to type this:(
1) Evaluate: lim 3x^4+4x^2-2/3x^3-5x^4 when x approaches infinity(the symbol looks like a sideways eight). I hope this makes sense because I need help.
2) If y=2x^2-3x,find the average rate of change of y with respect to x from P(2,2) to Q(2.1,2.52).

Found 2 solutions by josmiceli, Alan3354:
Answer by josmiceli(19441)

This is advanced algebra and you're not used to
seeing the infinity sign. I think you need lots
of basic reading, maybe tutoring. Infinity
should be like an old friend if you're going
to tackle this stuff - /: |

Answer by Alan3354(69443)

I am not even sure how to type this:(
1) Evaluate: lim 3x^4+4x^2-2/3x^3-5x^4 when x approaches infinity(the symbol looks like a sideways eight). I hope this makes sense because I need help.
Use L'Hopital's rule.
Limit = -0.6
email me via the thank you note if you need more info
-------------------
2) If y=2x^2-3x,find the average rate of change of y with respect to x from P(2,2) to Q(2.1,2.52)
For average, find f(2) and f(2.1)
f(2) = 2
f(2.1) = 8.82 - 6.3 = 2.52
Avg = diff in y/diff in x = 0.52/0.1 = 5.2
----------------------------
The slope of the function is 4x - 3.
At 2, m = 5
At 2.1, m = 5.4
The average is between the two slopes. It is the average of the 2 slopes in this case, but will not always be.